OpenStudy (anonymous):

Bill has 5 times as many baseball cards as Fred. Time has 3 fewer than 4 times as many baseball cards as Fred. In all they have 167 baseball cards. How many does each have?

OpenStudy (espex):

Fred=x \[5x+(4x-3)+x=167\]\[5x +x +4x-3=167 \rightarrow 10x=170 \rightarrow x=17\] So Fred has 17 cards, Bill has 5 times Fred, (5 x 17), and Tim has ((4 x 17)-3) Which means that: Bill has 85 cards Fred has 17 cards Tim has 65 cards 85+17+65 = 167